A 20.0 mL of 0.36 M solution of the salt NaA has a pH of 8.40.          

a) Calculate the pKa value of the acid HA. (Correct to 3 sig. fig)

b) Calculate the pH of a solution containing 0.4 M HA and 0.6 M NaA. (Correct to 3 sig. fig)
Answer #1

Here and pH = 8.40 pH = -log [4+] 8.40 = -log 4+ (ht=10-8.4 = 3-98410 -7M. [04-] = kw Kw = 1810-14 [ht] 1710-14 = 2.51x10-6 M(b) now [HA) = 0.4m (MaA] 20.0m 2 T RS NGA - Mat + A- so [A-] = 0.6M ANT Mob pu= pka + log [A-] 1. THA pH = 3.24 + log (0.6)